OpenStudy (anonymous):

Help please! Find the domain of the function: 5x+4/ x^2+3x+2

jimthompson5910 (jim_thompson5910):

you cannot divide by zero

jimthompson5910 (jim_thompson5910):

so if x^2+3x+2 were zero, then x^2+3x+2 = 0 (x+1)(x+2) = 0 x+1 = 0 or x+2 = 0 x = -1 or x = -2

jimthompson5910 (jim_thompson5910):

this means that if x = -1 or x = -2, then x^2+3x+2 = 0 so this tells us that the domain is the set of all real numbers but x cannot equal -1 and it cannot equal -2 (to avoid dividing by zero)

OpenStudy (anonymous):

Thanks!